1. Radiation machines, used to treat tumors, produce an inensity of radiation that valies inversely as the square of the distance from the machine. At 3 meters, the radiation intensity is 62.5 milliroentgens per hour. What is the intensity at a distance of 3.3 meters?
2. A moving sidewlak at an airport glides at a rate of 1.2 feet per second. Walking an the moving sidewalk, you travel 120 feet forward in the same time it takes to travel 24 feet in the oppsite direction. What is your walking speed on a nonmoving sidewalk?

Answer by edjones(8007)

Let x=radiation at the machine.
x/3^2
=x/9
x/9=62.5
x=9*62.5
=562.5 miliroentgens
.
562.5 * 1/3.3^2
=562.5/10.89
=51.65 miliroentgens
.
Ed